# FedAIoT: A Federated Learning Benchmark for Artificial Intelligence of Things

### v2

@@ 3.2.2 IoT-specific Data Preprocessing
The eight datasets included in FedAIoT cover diverse IoT data modalities such as wireless signals, drone images, and smart home sensor data. FedAIoT incorporates a suite of IoT-specific data preprocessing techniques that are designed for different IoT data modalities accordingly.
WISDM: We followed the standard preprocessing techniques used in accelerometer and gyroscope-based activity recognition for WISDM Ravi et al. (2005); Reyes-Ortiz et al. (2016); Ronao and Cho (2016). Specifically, for each $3$-minute session, we used a $10$-second sliding window with $50$% overlap to extract samples from the raw accelerometer and gyroscope data sequences. We then normalize each dimension of the extracted samples by removing the mean and scaling to unit variance.
@@
CASAS: We followed Liciotti et al. (2019) to transform the sensor readings into categorical sequences, creating a form of semantic encoding. Each unique temperature setting is assigned a distinct categorical value, as are individual instances of motion sensors and door sensors being activated. For each recorded activity, we then extract a sequence of the $2,000$ previous sensor activations which is used for activity modeling and prediction.
@@
EPIC-SOUNDS: We first performed a Short-Time Fourier Transform (STFT) on raw audio data followed by applying a Hanning window of 10ms duration and a step size of 5ms to ensure optimal spectral resolution. We then extracted $128$ Mel spectrogram features. To further enhance the data, we applied a natural logarithm scaling to the Mel spectrogram output and lastly, we padded each segment to reach a consistent length of $400$.
@@ 3.2.3 IoT-friendly Models
Since FedAIoT focuses on IoT devices which are resource-constrained, our choice of models is informed by a combination of model accuracies (Reinbothe, 2023; Yang et al., 2023; Terven and Cordova-Esparza, 2023; Liciotti et al., 2019; Seyedzadeh et al., 2018; Sholahudin et al., 2016; Huh et al., 2023) and model efficiency. For each included dataset, we evaluated multiple model candidates, and selected the best-performing ones that adhere to the resource constraints of representative IoT devices listed in Table 4. As an example, for UT-HAR dataset, two model candidates (ViT and ResNet18) have similar accuracy, and we chose the more efficient ResNet18 in our benchmark. Table 3 lists the selected model for each dataset. The detail of the architecture of each model is described in Appendix A.5.
@@ 3.2.4 FL Hyperparameters
FL Optimizer. FedAIoT supports a handful of commonly used FL optimizers. In the experiment section, we showcase the benchmark results of two of the most commonly used FL optimizers: FedAvg (McMahan et al., 2017) and FedOPT Reddi et al. (2020).
Client Sampling Ratio. Client sampling ratio denotes the proportion of clients selected for local training in each round of FL. This hyperparameter plays a crucial role as it directly influences the computation and communication costs associated with FL. FedAIoT facilitates the creation of diverse client sampling ratios and the evaluation of its impact on both model performance and convergence speed during FL training.
@@ 3.2.5 IoT-Factor Emulator
Noisy Labels. In real-world scenarios, IoT sensor data can be mistakenly labeled. Such data labeling errors act as noises in the federated training process. The concept of label noise has been discussed in Mazumder et al. (2024) for centralized training, which outlines several ways to select sub-samples of data to maximize data quality. However, in FL, this may not be practical as the data is private. Therefore, FL systems have to evaluate the robustness to label noise as an essential part of design. The impact of noisy labels on FL was first investigated in Zhang et al. (2023a). However, the label error distribution was considered uniform. In contrast, we propose a new probabilistic label error scheme based on the insight that labels that are similar to each other are more likely to be mislabeled. This design lets us simulate label errors in a more realistic manner. Specifically, we augment the ground truth labels of a dataset with a confusion matrix $Q$ where, $Q_{ij}$ is the probability that the true label $i$ is changed to a different label $j$, i.e., $P(\hat{y}=j\mid y=i)$. The confusion matrix was constructed based on centralized training results, where the elements of $Q_{ij}$ was determined from the ratio of samples labeled as $j$ by a centrally trained model to those with ground truth label $i$.
Quantized Training. IoT devices are resource-constrained. Previous benchmarks like FLUTE (Dimitriadis et al., 2022) examined the performance of quantized training during server-side model aggregation for the purpose of reducing the communication cost of FL. In comparison, FedAIoT incorporates not only quantized model aggregation at the server side but also quantized training at the client side to reduce the memory demands of FL on the device. Quantized training on both server and client sides is key to enabling FL for AIoT as IoT devices are not just restrained in communication bandwidth but also in on-device memory.
4 Experiments and Analysis
We implemented FedAIoT using PyTorch (Paszke et al., 2019) and Ray (Moritz et al., 2018), and conducted our experiments on NVIDIA A6000 GPUs. We ran each of our experiments on three random seeds and report the mean and standard deviation.
@@ 4.1 Overall Performance
Benchmark Results: Table 5 summarizes our results. We make three observations. (1) Data heterogeneity level and FL optimizer have different impacts on different datasets. In particular, the performance of UT-HAR, AEP, and EPIC-SOUNDS is very sensitive to the data heterogeneity level. In contrast, WISDM-P, CASAS, and VisDrone show limited accuracy differences under different data heterogeneity levels. (2) Under high data heterogeneity, FedAvg has a better performance compared to FedOPT in WISDM and Widar datasets and also has lower deviation for all datasets except WISDM-P and EPIC-SOUNDS. The performance gap between the two FL optimizers reduces under low data heterogeneity. (3) Compared to the other datasets, CASAS, AEP, and WISDM-W have higher accuracy margins between centralized training and low data heterogeneity.
@@ 4.4 Performance on Quantized Training
Lastly, we examine the impact of quantized training on FL under half-precision (FP16) [^1] PyTorch does not support lower quantization levels like INT8 and INT4 during training as referenced Team (2023) at the time of writing and hence those were excluded. Additionally, FP8 is only supported for H100 GPUs, not the mobile GPUs used for IoT devices. We assess model accuracy and memory usage under FP16 and compare the results to those from the full-precision (FP32) models. Memory usage is measured by analyzing the GPU memory usage of a model when trained with the same batch size under a centralized setting. Note that we use memory usage as the metric since it is a relatively consistent and hardware-independent metric. In contrast, other metrics such as computation speed and energy are highly hardware-dependent. Depending on the chipset that the IoT devices use, the computation speed and energy could exhibit wide variations. More importantly, new and more advanced chipsets are produced every year. The updates of the chipsets would inevitably make the benchmarking results quickly obsolete and out of date.
@@ 4.5 Insights from Benchmark Results
Need for Resilience on High Data Heterogeneity: As shown in Table 5, datasets can be sensitive to data heterogeneity. We observe that UT-HAR, Widar, AEP, and EPIC-SOUNDS show a significant impact under high data heterogeneity. These findings emphasize the need for developing advanced FL algorithms for data modalities that are sensitive to high data heterogeneity. Handling high data heterogeneity is still an open question in FL (Zhao et al., 2018; Li et al., 2020b) and our benchmark shows that it is a limiting factor for some of the IoT data modalities as well. To mitigate high data heterogeneity, although many techniques have been proposed (Sattler et al., 2019; Arivazhagan et al., 2019), their performance is still constrained. Most recently, some exploration on incorporating generative pre-trained transformers (GPT) as part of the FL framework has shown great performance in mitigating high data heterogeneity (Zhang et al., 2023b).
@@
Need for Resilience on Noisy Labels: As shown in Table 7, certain datasets exhibit high sensitivity to label errors, significantly deterring FL performance. Notably, both WISDM-W and Widar experience a drastic decrease in accuracy when faced with a $10$% label noise ratio. Given the inevitability of noise in real FL deployments where private data is unmonitored except by the respective data owners, the development of label noise resilient techniques becomes crucial for achieving reliable FL performance. However, handling noisy labels is a much less explored topic in FL. Pioneering work on this topic is still exploring the effects of noisy labels on the performance of FL (Zhang et al., 2023a; Feng et al., 2023). To mitigate the effect of noisy labels, for low label noise rates, techniques such as data augmentation and regularization have been shown to be effective in centralized training settings (Shorten and Khoshgoftaar, 2019; Zhang et al., 2017). If the label error rate is high, techniques such as knowledge distillation (Hinton et al., 2015), mixup Zhang et al. (2017), and bootstrapping (Reed et al., 2014) have been proposed. We expect such techniques or their variants would help in the FL setting.
Need for Quantized Training: Table 4 highlights the need for quantized training given the limited RAM resources on representative IoT devices. From our results summarized in Table 8, we observe that FP16 quantization does not affect accuracy drastically unless normalization layers are present like batch normalization (Jacob et al., 2018). To mitigate this issue, techniques such as incremental quantization (Zhou et al., 2017) or ternary weight networks (Li et al., 2016) have shown great performance in the centralized training setting. We expect such techniques or their variants would help in the FL setting as well.
@@ 5 Conclusion
In this paper, we presented FedAIoT, an FL benchmark for AIoT. FedAIoT includes eight datasets collected from a wide range of authentic IoT devices as well as a unified end-to-end FL framework for AIoT that covers the complete FL-for-AIoT pipeline. We have benchmarked the performance of the datasets and provided insights on the opportunities and challenges of FL for AIoT. Moving forward, we aim to foster community collaboration by launching an open-source repository for this benchmark, enriching it with additional datasets, algorithms, and analytical evaluations, to ensure it remains a valuable resource for FL for AIoT research.
Appendix A
A.1 Datasets used in FL Landscape
We summarize the datasets used by both seminal works and recent papers in federated learning in Table 9. A key reason for using these datasets is the availability of documentation and resources, allowing researchers to adopt them for testing quickly. However, these are not tailored for FL use cases, which highlights the need for good benchmarking tools for datasets that are more appropriate for FL scenarios as we did for AIoT in this work.
Table 9: Overview of the different datasets used in FL papers
@@ A.1 Datasets used in FL Landscape
width = colspec = Q[450]Q[781], column2 = c, hline1,15 = -0.08em, hline2 = -, Method & Dataset Used
FedAvg McMahan et al. (2017) MNIST LeCun et al. (1998), Shakespeare Karpathy (2015)
FedOPT Reddi et al. (2020) CIFAR-10/100 Krizhevsky et al. (2009), EMNIST Cohen et al. (2017), Shakespeare Karpathy (2015), StackOverflow TensorFlow Federated (2019)
Scaffold Karimireddy et al. (2020) EMNIST Cohen et al. (2017)
FedProx Li et al. (2020a) MNISTLeCun et al. (1998), FEMNIST Xiao et al. (2017), Shakespeare Karpathy (2015), Sent 140 Goel et al. (2016)
FedDF Lin et al. (2020) CIFAR-10/100 Krizhevsky et al. (2009), AG News Zhang et al. (2015), SST2 Socher et al. (2013)
DS-FL Itahara et al. (2021) MNIST LeCun et al. (1998), FEMNIST Xiao et al. (2017)
Fed-ETCho et al. (2022) CIFAR-10/100 Krizhevsky et al. (2009), Sent 140 Goel et al. (2016)
HeteroFL Diao et al. (2021) MINIST LeCun et al. (1998), CIFAR-10/100 Krizhevsky et al. (2009), WikiText Merity et al. (2017)
FedRolex Alam et al. (2022) CIFAR-10/100 Krizhevsky et al. (2009), StackOverflow TensorFlow Federated (2019)
FedGen Venkateswaran et al. (2023) SST-2 Socher et al. (2013), AG News Zhang et al. (2015), UT-HAR Yousefi et al. (2017), Water Quality Environmental Protection Agency (2021), BPIC 2018 van Dongen and Borchert (2018)
DynaFed Pi et al. (2023) FMNIST Xiao et al. (2017), CIFAR-10 Krizhevsky et al. (2009), CINIC-10 Darlow et al. (2018)
GPT-FL Zhang et al. (2023b) CIFAR-10/100 Krizhevsky et al. (2009) , Flower-102 Lv et al. (2021), Google Command Warden (2018), ESC-50 Piczak (2015)
pFedHR Wang et al. (2024) MNIST LeCun et al. (1998), SVHN Yang et al. (2021), and CIFAR-10 Krizhevsky et al. (2009)
A.2 Dirichlet Distribution
where $\Gamma(\cdot)$ is the gamma function.
A.3 Dataset Details

### v3

@@ 3.2.2 IoT-specific Data Preprocessing
The eight datasets included in FedAIoT cover diverse IoT data modalities such as wireless signals, drone images, and smart home sensor data. FedAIoT incorporates a suite of IoT-specific data preprocessing techniques that are designed for such IoT data modalities accordingly.
WISDM: We followed the standard preprocessing techniques used in accelerometer and gyroscope-based activity recognition for WISDM (Ravi et al., 2005; Reyes-Ortiz et al., 2016; Ronao and Cho, 2016). Specifically, for each $3$-minute session, we used a $10$-second sliding window with $50$% overlap to extract samples from the raw accelerometer and gyroscope data sequences. We then normalize each dimension of the extracted samples by removing the mean and scaling to unit variance.
@@
CASAS: We followed Liciotti et al. (2019) to transform the sensor readings into categorical sequences, creating a form of semantic encoding. Every possible temperature setting is assigned a distinct categorical value. Individual instances of motion and door sensor activation (on or off) are also assigned a categorical value. Subsequently, for a sensor activation, we extract a sequence consisting of the previous $2,000$ sensor activations which we then use for activity prediction.
@@
EPIC-SOUNDS: We first performed a Short-Time Fourier Transform (STFT) on raw audio data followed by applying a Hanning window of 10ms duration and a step size of 5ms to ensure appropriate spectral resolution. We then extracted $128$ Mel spectrogram features. To further enhance the data, we applied a natural logarithm scaling to the Mel spectrogram output. Lastly, we padded each segment to reach a consistent length of $400$.
@@ 3.2.3 IoT-friendly Models
Given that FedAIoT focuses on resource-constrained IoT devices, our choice of models is informed by a combination of model accuracies and model efficiency (Reinbothe, 2023; Yang et al., 2023; Terven et al., 2023; Liciotti et al., 2019; Seyedzadeh et al., 2018; Sholahudin et al., 2016; Huh et al., 2023). For each included dataset, we evaluated multiple model candidates, and selected the best-performing ones that adhere to the resource constraints of representative IoT devices listed in Table 4. As an example, for UT-HAR dataset, two model candidates (ViT and ResNet18) have similar accuracy, and we selected ResNet18 in our benchmark as it has better model efficiency. Table 3 lists the selected model for each dataset. The detail of the architecture of each model is described in Appendix D.
@@ 3.2.4 FL Hyperparameters
FL Optimizer. FedAIoT supports a handful of commonly used FL optimizers. In the experiment section, we showcase the benchmark results of two of the most commonly used FL optimizers: FedAvg (McMahan et al., 2017) and FedOPT (Reddi et al., 2021).
Client Sampling Ratio. Client sampling ratio denotes the proportion of clients selected for local training in each round of FL. This hyperparameter plays a crucial role as it directly influences the computation and communication costs associated with FL. FedAIoT facilitates the creation of diverse client sampling ratios and the evaluation of its impact on both model performance and convergence speed during federated training.
Details on other hyperparameters used for each experiment are described in Appendix E.
@@ 3.2.5 IoT-Factor Emulator
Noisy Labels. In real-world scenarios, IoT sensor data can be mistakenly labeled. These labels introduce noise into the federated training process. Therefore, FL systems need to be designed with robustness to label noise as a crucial feature. Label noise is a well-discussed concept in centralized training with established methods to mitigate it (Song et al., 2023). In general, label noise can be emulated in two ways: Symmetric Noise where the injected noise is uniformly distributed across the labels (Tanaka et al., 2018); and Asymmetric Noise injects noise non-uniformly across different labels using a label transition probability matrix (Patrini et al., 2017; Tanaka et al., 2018). In the context of FL, Xu et al. (2022); Zhang et al. (2023); Feng et al. (2023) propose to use symmetric noise to simulate label noise, evaluating FL algorithms at various noise levels. Fang and Ye (2022); Kim et al. (2022); Wu et al. (2023); Liang et al. (2024); Tsouvalas et al. (2024), on the other hand, propose to use asymmetric noise to simulate label noise. In particular, Fang and Ye (2022); Kim et al. (2022); Liang et al. (2024); Tsouvalas et al. (2024) design label transition probabilities in a pairwise manner, assuming that one label can only be mislabeled as another specific label with a random probability. In contrast, in FedAIoT, we propose a new label transition probability matrix that breaks this assumption, where a label can be mistakenly labeled as any other label with a learned probability. Specifically, we augment the ground truth labels of a dataset with a label transition probability matrix $Q$ where, $Q_{ij}$ is the probability that the true label $i$ is changed to a different label $j$, i.e., $P(\hat{y}=j\mid y=i)$. The label transition probability matrix was constructed based on centralized training results, where the elements of $Q_{ij}$ was determined from the ratio of samples labeled as $j$ by a centrally trained model to those with ground truth label $i$.
Quantized Training. IoT devices are resource-constrained. Previous benchmarks such as FLUTE (Garcia et al., 2022) examined the performance of quantized training during server-side model aggregation to reduce the communication cost of FL. Ozkara et al. (2021) and Gupta et al. (2023) study quantization-aware training but still use full precision during training. In contrast, FedAIoT incorporates not only quantized model aggregation at the server side but also quantized training at the client side to reduce the memory demands of FL on client IoT devices. Quantized training on both server and client sides is key to enabling FL for AIoT as IoT devices are not just restrained in communication bandwidth but also in on-device memory.
4 Benchmark Results and Analysis
We implemented FedAIoT using PyTorch (Paszke et al., 2019) and Ray (Moritz et al., 2018) and conducted our experiments on NVIDIA A6000 GPUs. We run each of our experiments using three random seeds and report the mean and standard deviation.
@@ 4.1 Overall Performance
Benchmark Results: Table 5 summarizes our results. We make three observations. (1) Data heterogeneity level and FL optimizer have different impacts on different datasets. In particular, the performance of UT-HAR, AEP, and EPIC-SOUNDS is extremely sensitive to the data heterogeneity level. In contrast, WISDM-P, CASAS, and VisDrone show limited accuracy differences under different data heterogeneity levels. (2) Under high data heterogeneity, FedAvg has a better performance compared to FedOPT in WISDM and Widar datasets and has lower deviation for all datasets except WISDM-P and EPIC-SOUNDS. The performance gap between the two FL optimizers reduces under low data heterogeneity. (3) Compared to the other datasets, CASAS, AEP, and WISDM-W have higher accuracy margins between centralized training and low data heterogeneity.
@@ 4.4 Performance on Quantized Training
Lastly, we examine the impact of quantized training on FL under half-precision (FP16) [^1] PyTorch does not support lower quantization levels like INT8 and INT4 during training as referenced Pytorch (2023) at the time of writing and hence those were excluded. Additionally, FP8 is only supported for H100 GPUs, not the mobile GPUs used for IoT devices.. We assess model accuracy and memory usage under FP16 and compare the results to those from the full-precision (FP32) models. Memory usage is measured by analyzing the GPU memory usage of a model when trained with the same batch size under a centralized setting. Note that we use memory usage as the metric since it is a relatively consistent and hardware-independent metric. In contrast, other metrics such as computation speed and energy are highly hardware-dependent. Depending on the chipset that the IoT devices use, the computation speed and energy could exhibit wide variations. More importantly, new and more advanced chipsets are produced every year. The updates of the chipsets would inevitably make the benchmarking results quickly obsolete and out of date.
@@ 4.5 Insights from Benchmark Results
Need for Resilience on High Data Heterogeneity: As shown in Table 5, datasets can be sensitive to data heterogeneity. We observe that UT-HAR, Widar, AEP, and EPIC-SOUNDS show a significant impact under high data heterogeneity. These findings emphasize the need for developing advanced FL algorithms for data modalities that are sensitive to high data heterogeneity. Handling high data heterogeneity is still an open question in FL (Zhao et al., 2018; Li et al., 2020a) and our benchmark shows that it is a limiting factor for some of the IoT data modalities as well. To mitigate high data heterogeneity, although many techniques have been proposed (Sattler et al., 2019; Arivazhagan et al., 2019), their performance is still constrained. Most recently, some exploration on incorporating generative pre-trained transformers (GPT) as part of the FL framework has shown great performance in mitigating high data heterogeneity (Zhang et al., 2024).
@@
Need for Resilience on Noisy Labels: As shown in Table 7, certain datasets exhibit high sensitivity to label errors, significantly deterring FL performance. Notably, both WISDM-W and Widar experience a drastic decrease in accuracy when faced with a $10$% label noise ratio. Given the inevitability of noise in real FL deployments where private data is unmonitored except by the respective data owners, the development of label noise resilient techniques becomes crucial for achieving reliable FL performance. However, handling noisy labels is a much less explored topic in FL. Pioneering work on this topic is still exploring the effects of noisy labels on the performance of FL (Zhang et al., 2023; Feng et al., 2023). To mitigate the effect of noisy labels, for low label noise rates, techniques such as data augmentation and regularization have been shown to be effective in centralized training settings (Shorten and Khoshgoftaar, 2019; Zhang et al., 2018). If the label error rate is high, techniques such as knowledge distillation (Hinton et al., 2015), mixup (Zhang et al., 2018), and bootstrapping (Reed et al., 2015) have been proposed. We expect such techniques or their variants would help in the FL setting.
Need for Quantized Training: Table 4 highlights the need for quantized training given the limited RAM resources on representative IoT devices. From our results summarized in Table 8, we observe that FP16 quantization does not affect accuracy drastically unless normalization layers are present, like batch normalization (Jacob et al., 2018). To mitigate this issue, techniques such as incremental quantization (Zhou et al., 2017) or ternary weight networks (Liu et al., 2023) have shown great performance in the centralized training setting. We expect such techniques or their variants would help in the FL setting as well.
@@ 5 Conclusion
In this paper, we presented FedAIoT, an FL benchmark for AIoT. FedAIoT includes eight datasets collected from a wide range of authentic IoT devices as well as a unified end-to-end FL framework for AIoT that covers the complete FL-for-AIoT pipeline. We have benchmarked the performance of the datasets and provided insights on the opportunities and challenges of FL for AIoT. Moving forward, we aim to foster community collaboration by launching an open-source repository for this benchmark, continually enriching it with more datasets, algorithms, and thorough analytical evaluations, ensuring it remains a dynamic and invaluable resource for FL for AIoT research.
6 Acknowledgement
We would like to thank the action editor Peter Mattson and the anonymous reviewers of the Journal of Data-centric Machine Learning Research for their helpful and constructive comments. Zhichao Cao is supported in part by National Science Foundation (NSF) under award NeTS-2312675. JeongGil Ko is partially supported by the Ministry of Science and ICT (MSIT) and the IITP (IITP-2022-2020-001461, IITP-2022-0-00420) in the Republic of Korea. Tiantian Feng and Shrikanth S. Narayanan are supported in part by USC + Amazon Center on Secure & Trusted ML. Tuo Zhang and Salman Avestimehr are supported in part by ARO award W911NF1810400, ONR Award No. N00014-16-1-2189 and NSF under NSF SATC ProperData Center. Samiul Alam and Mi Zhang are supported in part by the Meta Reality Labs Faculty Research Award. The views, opinions, and/or findings expressed are those of the author(s) and should not be interpreted as representing the official views or policies of the Department of Defense or the U.S. Government.
Appendix A Datasets used in Existing Representative FL Works
Table 9: Summary of the datasets used in existing representative FL works.
@@ Appendix A Datasets used in Existing Representative FL Works
width = colspec = Q[450]Q[781], column2 = l, hline1,15 = -0.08em, hline2 = -,                  Method &                          Datasets Used
FedAvg (McMahan et al., 2017) MNIST (LeCun et al., 1998), Shakespeare (Karpathy, 2015)
FedOPT (Reddi et al., 2021) CIFAR-10/100 (Krizhevsky et al., 2009), EMNIST (Cohen et al., 2017), Shakespeare (Karpathy, 2015), StackOverflow (TensorFlow Federated, 2019)
Scaffold (Karimireddy et al., 2020) EMNIST (Cohen et al., 2017)
FedProx (Li et al., 2020b) MNIST(LeCun et al., 1998), FEMNIST (Xiao et al., 2017), Shakespeare (Karpathy, 2015), Sent 140 (Goel et al., 2016)
FedDF (Lin et al., 2020) CIFAR-10/100 (Krizhevsky et al., 2009), AG News (Zhang et al., 2015), SST2 (Socher et al., 2013)
DS-FL (Itahara et al., 2021) MNIST (LeCun et al., 1998), FEMNIST (Xiao et al., 2017)
Fed-ET(Cho et al., 2022) CIFAR-10/100 (Krizhevsky et al., 2009), Sent 140 (Goel et al., 2016)
HeteroFL (Diao et al., 2021) MINIST (LeCun et al., 1998), CIFAR-10/100 (Krizhevsky et al., 2009), WikiText (Merity et al., 2017)
FedRolex (Alam et al., 2022) CIFAR-10/100 (Krizhevsky et al., 2009), StackOverflow (TensorFlow Federated, 2019)
FedGen (Venkateswaran et al., 2023) SST-2 (Socher et al., 2013), AG News (Zhang et al., 2015), UT-HAR (Yousefi et al., 2017), Water Quality (Environmental Protection Agency, 2021), BPIC 2018 (van Dongen and Borchert, 2018)
DynaFed (Pi et al., 2023) FMNIST (Xiao et al., 2017), CIFAR-10 (Krizhevsky et al., 2009), CINIC-10 (Darlow et al., 2018)
GPT-FL (Zhang et al., 2024) CIFAR-10/100 (Krizhevsky et al., 2009) , Flower-102 (Lv et al., 2021), Google Command (Warden, 2018), ESC-50 (Piczak, 2015)
pFedHR (Wang et al., 2024) MNIST (LeCun et al., 1998), SVHN (Yang et al., 2021), and CIFAR-10 (Krizhevsky et al., 2009)
Appendix B Dataset Details
B.1 WISDM
The WISDM dataset comprises raw accelerometer and gyroscope data collected from 51 subjects performing 18 activities for three minutes each. Data were gathered at a 20Hz sampling rate from both a smartphone (Google Nexus 5/5x or Samsung Galaxy S5) and a smartwatch (LG G Watch). Data for each device and sensor type are stored in different directories, resulting in four directories overall. Each directory contains 51 files, each corresponding to a subject. The data entry format is: <subject-id, activity-code, timestamp, x, y, z>. Separate files for the gyroscope and accelerometer readings are provided and are later combined by matching timestamps. Subject ID is given from 1600 to 1650 and the activity code is an alphabetical character between ‘A’ and ‘S’ excluding ‘N’. The timestamp is in Unix time. The code to read and partition the data into 10s segments is provided by our benchmark. The input shape of the processed data is $200\times 6$. The original dataset is available at https://archive.ics.uci.edu/dataset/507/wisdm+smartphone+and+smartwatch+activity+and+biometrics+dataset.
B.2 UT-HAR
B.3 Widar
B.4 VisDrone
B.5 CASAS
The CASAS dataset is a collection of data generated in smart home environments, where intelligent software uses sensors deployed at homes to monitor resident activities and conditions within the space. The CASAS project considers environments as intelligent agents and employs custom IoT hardware known as Smart Home in a Box (SHiB), which encompasses the necessary sensors, devices, and software. The sensors in SHiB perceive the status of residents and their surroundings, and through controllers, the system acts to enhance living conditions by optimizing comfort, safety, and productivity. The CASAS dataset includes the date (in yyyy-mm-dd format), time (in hh:mm:ss.ms format), sensor name, sensor readings, and an activity label in string format. The data were collected in real-time as residents go about their daily activities. The code to extract categorical sensor readings to create input sequences and labels is provided in our benchmark. The CASAS dataset can be downloaded from https://casas.wsu.edu/datasets/.
B.6 AEP
B.7 EPIC-SOUNDS
Table 10: Details of the datasets included in FedAIoT.